Using the add-on method, calculate the APR of a loan for $11,650 at 10.5% for 2 years. (Do not round intermediate calculations. Round your answer to the nearest tenth percent.)
Answer #1

We know that APR = loan amount * interest rate/ term of loan

Thus APR =11650 * 10.50/100/ 48= Rs 1223.25/24 = Rs 50.96875 per month

wherein loan amount = Rs 11650

Interest rate = 10.50%

term of loan = 24 months